# A High Efficient and Scalable Obstacle-Avoiding VLSI Global Routing Flow

### v2

@@ 6.4 Proposed Obstacle-Avoiding Global Routing Evaluations
6.5 Obstacle-Avoiding Global Routing Insight Analysis
Besides the results of the complete flow, we also analyze the insides, including runtime breakdown, the comparisons for initial routing step, the comparisons of using different graphs during maze routing, and the effectiveness of OARSMT-guided sparse maze routing
@@ 6.5 Obstacle-Avoiding Global Routing Insight Analysis
Next, Figure 10 illustrates the obstacle violations and runtime ratio of initial routing using the advanced OARSMT (“Lin et al.”) and our initial routing (“Ours”) to CUGR 2.0. Our initial routing has from 70% to 95% fewer obstacle violation cost than CUGR 2.0, from 16% to 71% fewer than “Lin et al.”. “Lin et al.” is unable to generate effective solutions in some cases, which results in a higher obstacle violation cost during the initial routing phase compared to “Ours”. Furthermore, “Ours” has $\sim$28x - 177x speedup compared with “Lin et al.”, with 149x on average. This again indicates the advantages of the proposed OARSMT algorithm on quality and speed towards complete designs.
Figure 10: Obstacle violation and runtime ratio of the initial routing using the Lin et al. OARSMT and our initial routing to the CUGR 2.0. (The testcases with and without “*” are same meaning as those in Table 3.)
@@
Then, the comparisons between different rip-up and reroute approaches are analyzed. Table 4 shows the normalized metrics of conventional maze routing with dense routing graphs, compared to the proposed rip-up and reroute step, which includes OARSMT-guided sparse maze routing and obstacle-aware sparse maze routing. The initial routing steps both apply the proposed algorithms. It can be noticed that, although using dense routing graph in conventional maze routing can have better wirelength and overflow, it results in an average 71x increase in running time. For the second largest benchmark mempool_group, the proposed method requires approximately only one hour of runtime, whereas using conventional maze routing takes over 3 days. For the largest case clauster, the conventional maze routing cannot be finished in 4 days. Due to the large runtime particularly for complex cases, using dense graph for maze routing is impractical and the proposed approach is efficient.
@@
Furthermore, we analyzed the effectiveness of OARSMT-guided sparse maze routing on obstacle violations and overflow. As shown in Figure 11, it illustrates the obstacle violations and overflow ratio of our obstacle-avoiding global routing flow without the obstacle-aware sparse maze routing to the enhanced CUGR 2.0. Proposed OARSMT-guided sparse maze routing can completely eliminate obstacle violations in 4 cases. For the remaining 3 larger cases, although they are also almost cleared, which is significantly less than the enhanced CUGR 2.0. In addition to mempool_tile, our OARSMT-guided sparse maze routing outperforms the enhanced CUGR2.0 in all other cases, with the best improvement reaching 58.7% and the average improvement being 15.3%. The absolute value of the overflow cost for mempool_tile is only 13 less than the proposed flow. This proves that using the proposed OARSMT algorithm to consider the obstacles in the early stage, and providing the guidance for maze routing, the efficiency of avoiding the obstacles in for rip-up and reroute can be improved.
@@ 7 Conclusion
This work proposes an effective obstacle-avoiding global routing. Leveraging the proposed rule-based OARSMT algorithm in initial routing, it significantly reduces the obstacle violations in the early stage. By proposing the OARSMT-guided sparse maze routing and obstacle-aware sparse maze routing in rip-up and reroute, the obstacle violations are effectively eliminated. Compared with the advanced OARSMT algorithm, traded with limited wirelength overhead, the proposed OARSMT algorithm has $\sim$10x-2700x and $\sim$150x-5800x runtime speedup under randomized testcases and standard benchmarks, respectively. Compared with the enhanced advanced global router CUGR 2.0, proposed obstacle-avoiding global router successfully eliminates obstacle violations, and reduces wirelength and overflow cost on the modified ISPD 2024 benchmarks, traded limited via counts and runtime overhead. Besides, it achieves an average improvement of 1.96% in wirelength and a 28.06% reduction in overflow cost.

### v3

@@ 6.5 Obstacle-Avoiding Global Routing Insight Analysis
6.5.1 Runtime Breakdown
@@
6.5.2 Comparisons of Different OARSMT Algorithms for Initial Routing
Next, Figure 12 illustrates the obstacle violations and runtime ratio of initial routing using the previously advanced OARSMT (“Lin et al.”) and our initial routing (“Ours”) to CUGR 2.0. Our initial routing has from 70% to 95% fewer obstacle violations than CUGR 2.0, from 16% to 71% fewer than “Lin et al.”. “Lin et al.” is unable to generate effective solutions in some cases, which results in a higher obstacle violation cost during the initial routing phase compared to “Ours”. Furthermore, “Ours” has $\sim$28x - 177x speedup compared with “Lin et al.”, with an average of 149x. This again indicates the advantages of the proposed OARSMT algorithm on quality and speed towards complete designs.
Figure 12: Obstacle violation and runtime ratio of the initial routing using the Lin et al. OARSMT and our initial routing to the CUGR 2.0.
@@ 6.5.2 Comparisons of Different OARSMT Algorithms for Initial Routing
6.5.3 Comparisons of Using Different Graphs during Maze Routing
@@
Then, the comparisons between different rip-up and reroute approaches are analyzed. Table 4 shows the normalized metrics for conventional maze routing with dense routing graphs, compared to the proposed rip-up and reroute step, which includes OARSMT-guided sparse maze routing and obstacle-aware sparse maze routing. The initial routing steps both apply the proposed algorithms. It can be noticed that, although using dense routing graphs in conventional maze routing can have better wirelength and overflow, it results in an average 71x increase in running time. For the second-largest benchmark mempool_group, the proposed method requires approximately only one hour of runtime, whereas using conventional maze routing takes over 3 days. For the largest case clauster, the conventional maze routing cannot be finished in 4 days.
Due to the large runtime, particularly for complex cases, using dense graphs for maze routing is impractical, and the proposed approach is efficient.
6.5.4 Effectiveness Analysis of OARSMT-Guided Sparse Maze Routing
@@
Furthermore, we analyzed the effectiveness of OARSMT-guided sparse maze routing in obstacle violations and overflow. As shown in Figure 13, it illustrates obstacle violations and the overflow ratio of our obstacle-avoiding global routing flow without obstacle-aware sparse maze routing compared to the enhanced CUGR 2.0. The proposed OARSMT-guided sparse maze routing can completely eliminate obstacle violations in 4 cases. For the remaining 3 larger cases, although they are also almost cleared, their violations are significantly less than the enhanced CUGR 2.0. Except for mempool_tile, our OARSMT-guided sparse maze routing outperforms the enhanced CUGR2.0 in all other cases, with the best improvement reaching 58.7% and the average improvement being 15.3%. This shows that using the proposed OARSMT algorithm to consider obstacles in the early stage and provide guidance for maze routing, the efficiency of avoiding the obstacles in for rip-up and reroute can be improved.
6.5.5 Ablation Experiments
Finally, we conduct an ablation study by isolating the rule-based OARSMT and our proposed maze routing, respectively. We calculated the average values of five metrics across all testcases for three methods: wirelength, via count, overflow, runtime, and obstacle violations. All metrics except obstacle violations are normalized ratios relative to our method. As shown in Figure 14, the performance of using only the rule-based OARSMT method is worse than “Ours”. Although the proposed OARSMT algorithm can handle most obstacles during tree generation, without the collaboration of the proposed maze routing, conventional sparse maze routing cannot leverage the guidance from our OARSMT structures, and thus, it cannot well handle the OARSMTs by its routing graph, leaving residual violations. On the other hand, only using the proposed maze routing without OARSMT generation also failed to provide the best results. As there is no guidance from OARSMT in the early stages, the obstacles are intensively considered during the maze routing step, where there are fewer opportunities to adjust the routing compared with the early steps. This results in higher overflow. This shows that both the OARSMT and the proposed maze routing are indispensable components, and applying them together can efficiently and synergistically handle the obstacles without scarifying much wirelength and via count.
Figure 14: Ablation study on routing performance.
6.6 Discussions of Optimality and Legality of the Proposed OARSMT Algorithm
In this section, we discuss the optimality and legality of the proposed OARSMT algorithm in Section 4.3 based on the experimental results.
For optimality, the proposed OARSMT algorithm is designed to support efficient obstacle-avoiding routing in practical scenarios. As the proposed algorithm spends polynomial time, it does not guarantee optimal solutions under the NP-hard nature of the OARSMT problem. However, the proposed algorithm is still able to provide slightly better solutions compared to previous works with higher efficiency, which is important for modern large-scale designs.
For legality, proving the proposed OARSMT algorithm can provide legal solutions for all cases mathematically is challenging. However, extensive experimental results (Tables 1 and 6.3) show that the proposed algorithm successfully avoids obstacles for all nets in the tested benchmarks, including the industrial testcases. This proves the effectiveness in practice. Even though there might be some corner cases in theory, where the OARSMT algorithm fails to provide a valid solution, its probability is quite limited (and it does not actually happen in the experiments). Moreover, these corner cases can also be resolved by rip-up and reroute step in the overall global routing flow. Note that there are obstacles violations during the initial routing step (such as Figures 10 and  12) in the proposed global routing flow, but these are different from the discussed theoretical corner cases. This is mainly due to the fact that when generating an OARSMT for a net during the initial routing, only obstacles within the net’s bounding box are considered to improve the efficiency, according to Section 5.1.
@@ 7 Conclusion
This work proposes an effective obstacle-avoiding global routing flow. Leveraging the proposed rule-based OARSMT algorithm in initial routing, it significantly reduces the obstacle violations in the early stage. By proposing the OARSMT-guided sparse maze routing and obstacle-aware sparse maze routing in rip-up and reroute, the obstacle violations are effectively eliminated. Compared with the previously advanced OARSMT algorithm, with limited wirelength overhead, the proposed OARSMT algorithm achieves $\sim$10x–2700x and $\sim$150x–5800x runtime speedup on randomized testcases and standard benchmarks, respectively. Compared with the previously advanced enhanced global router CUGR 2.0, the proposed obstacle-avoiding global router successfully eliminates obstacle violations, and reduces wirelength and overflow cost on the modified ISPD 2024 benchmarks, with limited via count and runtime overhead. In addition, it achieves an average improvement of 1.96% in wirelength and a reduction of 28.06% in overflow cost.